I’ll give thumbs up right away!
DONT HAVE TO SHOW WORK
1) if you invest 2000 each year for 40 years, how much will you have at the end of 40 years assuming 9% annual return?

2) if you want to withdraw 35000 each year for 30 years, how much would you need to have saved today to fund these withdrawals?

3) if you purchase a machine and borrow 10000, what will your approximate monthly payment be if you pay back the lon over 5 years? assume 10% interest rate.

4) you are considering purchasing a new machine for 112000 that will save 21000 per year for 8 years. if coat of capital is 12%, what is the net present value (NPV) of buying new machine?

5) you consider investing in a project that will require an initial capital outlay of 140000 and generate annual cash flow of 21805 for 10 years. what is the internal rate of return (IRR) for this project?

6) if you borrow 700000 to purchase a home, what do you expect the monthly payment will be? assume a 30 year mortgage and a fixed 7% interest rate.

7) if you invest 105000 each year for 35 years, how much will you have at the end of 35 years assuming 10% annual return?

8) you are considering investing 35000 in a project that will generate cash flows of 6000 per year for 12 years. if your cost of capital is 10%, what is the net present value of the investment?
Answer #1

Answer to Question 1:

Annual Investment = 2,000
Annual Interest Rate = 9%
Period = 40 years

Future Value = 2,000*1.09^39 + 2,000*1.09^38 + …. + 2,000*1.09 + 2,000
Future Value = 2,000 * (1.09^40 - 1) / 0.09
Future Value = 2,000 * 337.88245
Future Value = 675,764.90